• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 13:19:52 on November12,2002

    Entry number 89969

    keyword=Beam tuning

    Around 9:20 a.m. beam current was increased up to about 100 uA
    although it was not stable. MCC was supposed to take back the beam for
    tuning with new optic setting. They did not do the beam tuning till
    around noon. We were told that it might take two hours fo do the
    tuning.